OpenStudy (anonymous):

Richard's estate is to be split equally among his three children. If his estate is worth $1 3/4 million, how much will each children inherit?

OpenStudy (anonymous):

583,333 each with a dollar left over. I should get it rather than a medal.

OpenStudy (anonymous):

1 and 3/4 millions = 7/4 millions = 1.75 millions = 1750000 each child gets 1750000/3=583,333.33 dollars